# Inferring State Machine from the Protocol Implementation via Large Language Model

### v3

@@ III-C Context Window of LLM
Motivated by the findings discussed in the Section III, we propose an approach (ProtocolGPT) for inferring state machines based on LLM. The core idea of our approach is to harness the code analysis capabilities of large language models to infer protocol finite state machines from complex protocol implementation. To overcome the challenges aforementioned, we use the augmented LLM that combines the strengths of both retrieval models and generative models. Figure 2 illustrates the overall workflow of  ProtocolGPT, which consists of two primary components: LLM augmentation with source code and state machine inference.
@@
LLM Augmentation with Source Code. Initially, we isolate sections pertinent to the protocol state machine from the extensive and complex protocol implementation codebase. To guarantee efficient statement inference, the source code is segmented into smaller chunks to bypass limitations of LLM, such as the constraint on context window. Moreover, ProtocolGPT maintains the integrity of the original code structure by syntax-aware segmenting for classes, functions, and other structural components. Our method preserves the semantic and contextual information, mitigating potential losses by traditional text segmentation technique. By embedding the LLM with these code chunks,  ProtocolGPT could more precisely retrieve the most relevant code from vector store during the state machine inference.
State Machine Inference. To enable the augmented LLM to infer FSM, we integrate the chain-of-thought [23] and background-augmented prompting [51]. Leveraging domain-specific expertise, we formulate prompts tailored to the task at hand, which guide the augmented LLM in retrieving code segments that define both messages and states. The augmented LLM subsequently generates the message types and states based on the retrieved code. Finally, using the extracted message types and states, we traverse the entire state space to infer the corresponding transitions.
@@ IV-A LLM Augmentation with Source Code
While LLMs exhibit remarkable proficiency in code analysis, their efficacy diminishes significantly when tasked with analyzing large-scale, structurally complex programs. To address this limitation, ProtocolGPT leverages retrieval-augmented generation (RAG) techniques[52], which empower LLMs by integrating customizable external knowledge sources, such as code repositories. Although RAG models mitigate the inherent limitation of relying exclusively on the training data and learned patterns, their performance deteriorates significantly if the code repositories are not adequately preprocessed. Specifically, protocol implementations frequently include code snippets that exceed the LLM’s context window and often contain substantial amounts of code that are irrelevant to state machines, which may adversely impact retrieval accuracy, thereby compounding the challenges faced in effective code analysis. Consequently, preprocessing the code becomes a critical step in ensuring the LLM can be effectively augmented and leveraged for state machine inference.
First, the protocol implementation must be filtered to extract sections relevant to the state machine. The source code is then segmented into chunks to comply with the limitation of context window. The code chunks are then embedded and stored in a vector database, contributing to the enhancement of the pre-trained model’s performance by enabling efficient retrieval and preserving contextual relevance. During the inference phase, relevant code chunks are retrieved from the vector database and fed into the LLM, supporting the generation of the state machine.
@@ IV-A1 Code Filtering
To ensure scalability and maintainability, protocol implementations are typically structured into distinct modules, including state machines, configuration management, security and testing. Filtering out irrelevant code not only reduces the cost of LLM prompts but also enhances the accuracy and efficiency of FSM inference. Consequently, it is essential to focus on state machine module when inferring FSM from protocol implementations, as irrelevant modules do not contribute to the inference process. As discussed in Section III-C, the tokens in protocol implementations significantly exceeds the context window of LLM, preventing analysis of the implementation. Therefore, the first step in augmenting the LLM is to filter out the code related to the protocol state machine from the entire codebase.
State machine Components Identification. We observe that variables, constants, structures, and functions related to the state machine are often named according to terms defined in the RFCs. For instance, the CLIENT_HELLO message in s2n-tls[46] corresponds to the Client Hello message in the RFC. Additionally, keywords such as state and message frequently appear in these context-specific names. Code filter identifies the directory containing the state machine components within the protocol implementation through regular expression matching. Based on these observation, we define a comprehensive set of keywords related to protocol state machines, derived from both RFCs and expert knowledge. The keyword set comprises the message types and states as defined in the relevant RFC, supplemented by custom keywords based on expert knowledge. For instance, in the IKE protocol, Security Association (SA) appears frequently since the state of the SA essentially represents the state of the peer. These specially keywords facilitate the identification of state machine-related code. Furthermore, our keyword set can be seamlessly extended to accommodate additional protocols with negligible engineering effort, as these keywords are commonly used as section headers.
This pre-constructed keyword set is employed to perform regular expression matching on each document within the protocol implementation, and Matching documents are marked as state machine-related code. And the subdirectory where the documents have the highest hit rate will be selected as the state machine module, providing domain knowledge for LLM augmentation.
@@ IV-A2 Syntax-aware Code Segmentation
For the RAG model, the retrieved data is directly provided as context to the prompt, enabling the LLM to generate a response. To retrieve data from the vector database, the documents are transformed into chunks and embeded by the embedding model which also has a limited context window. However, traditional character-based splitting strategies risk disrupting semantically meaningful elements and may result in the loss of contextual information. Consequently, two critical factors must be taken into account during code segmentation: the size of code chunks and the syntactic structure of the code.
@@
Appropriate Chunk Size. Embeddings compress token-level representations into a single fixed-length vector, encapsulating the meaning of the entire input sequence. Appropriately sized chunks contain more precise contextual information, resulting in improved matching accuracy. The input provided to the LLM includes not only the retrieved documents but also instructions and relevant domain knowledge. However, if the length of the retrieved document approaches or exceeds the context window of LLM, the performance of the model may be adversely affected.
@@
Syntactic Structure of Program Language. Conventional text segmentation overlooks the structural composition of the document. When applied to data such as code, which is characterized by inherent syntax, these strategies can compromise essential components, including structural and functional elements. Moreover, the rich contextual information embedded within the code may be degraded during the segmentation process.
Figure 3: The workflow of Syntax-aware Code Segmentation Strategy.
@@
Building upon the preceding observations, we propose a syntax-aware code splitting strategy. This method break down protocol implementation documents into appropriately sized chunks, ensuring that the syntactic integrity of the code remains intact while also preserving a sufficient level of contextual information. Figure 3 illustrates the workflow of our proposed syntax-aware splitting technique. (1) Initially, we define a maximum allowable chunk size, referred to as MaxChunkSize. We then recursively decompose the documents using a series of language-specific separators until the size of each chunk is smaller than MaxChunkSize. These separators consist of keywords that define critical elements such as structures, classes, functions, and other constructs. For instance, in the C programming language, separators include terms such as struct, class, enum, static, etc. To facilitate the application of our method across multiple programming languages, we compiled a comprehensive list of separators for languages including C, C++, Python, and others. (2) To prevent the creation of overly small chunks that may lack sufficient contextual information, we establish a predefined minimum threshold for chunk size, named MinChunkSize. If the length of the chunk falls below this threshold, it is merged with an adjacent chunk to ensure coherence. The value of MinChunkSize is specified by users. (3) Finally, to preserve some level of contextual continuity between chunks, we regenerate new chunks based on the previously segmented ones, ensuring a specified degree of overlap is maintained between them.
@@ IV-A3 Vector Store
The augmented model employs semantic search to retrieve relevant documents from a vector store. Code snippets extracted from these documents are then appended to the original prompt, which is subsequently input into the generative model. Data retrieval from the vector store is facilitated by converting code into embeddings (high-dimensional vectors) that facilitate efficient search via a retrieval engine. The vector store is specifically optimized for storing embeddings of repository code from the protocol implementation, enabling the use of advanced search algorithm to identify matches between similar vectors.
When ProtocolGPT inferring state machine , it initially embeds the prompt to create an embedded version. It then retrieves the vectors from the vector store that most closely match the embedded prompt. Finally, the augmented LLM generate a more contextually relevant response based on the prompt posed by the original version and the code corresponding to these highly similar vectors.
For the storage and retrieval of unstructured code data, the optimal approach is to embed the data and store the resulting embedding vectors. Our approach leverages an embedding model provided by OpenAI[53] to transform code chunks into vector data. This vector data is subsequently stored and retrieved using a vector store facilitated by FAISS[54]. Each vector in the vector store corresponds to a specific piece of code data. The retrieval model employs the approximate nearest neighbor algorithm to retrieve the vectors relevant to the given query from the vector store.
@@ IV-B FSM Inference
Before inferring FSM using the augmented LLM, it is essential to define the FSM structure that the model generates. Our primary objective is to examine the transition relationships between different states within the protocol implementation. Consequently, we anticipate that the FSM generated by the LLM serves as a high-level abstract model that encompasses both the set of states within the protocol implementation and the transitions between these states. Specifically, the FSM details the potential states the protocol server can attain, the pairs of source and destination states, and the conditions necessary for transitioning from the source state to the destination state. To facilitate more effective support for downstream applications, FSM discussed in this paper is assumed to be a non-deterministic FSM that can be transformed into a deterministic FSM. Notably, an non-deterministic FSM may possess multiple initial states, and a single state can transition to several successor states in response to diverse inputs. Inspired by RFCNLP[9] and StateLifter[10], we define an non-deterministic FSM in Definition 1 as a formal model.
@@
Moreover, a significant challenge in utilizing augmented LLMs for downstream tasks is ensuring that their outputs are machine-readable. Unlike static and dynamic analysis methods, the inputs and outputs of generative model are formulated in natural language. Therefore, without imposing constraints on the output of model, the resulting FSM would be unusable. This challenge can be addressed through fine-tuning prompts, which guide the LLM to perform specific tasks and generate outputs according to the given patterns. This approach ensures that the output generated by the LLM adheres to our desired format.
The FSM inference phase leverages prompt engineering to direct the augmented LLM in deriving the FSM from the source code. However, this process is not instantaneous. As the complexity of a task increases with the number of steps, the probability of errors in the LLM’s output also rises. In contrast, the model demonstrates exceptional accuracy when executing single instructions. To mitigate this limitation, we employ the chain-of-thought technique [23], guiding the augmented LLM through each step of the process. Specifically, the FSM inference is structured into three distinct stages: (1) Retrieving the code snippets that define states, messages, and state transition relationships. (2) Extracting all the messages and states defined in the protocol implementation. (3) Identifying all the state transitions and the corresponding message types that trigger these transitions. By following these steps,  ProtocolGPT gradually extracts the information which forms the FSM of the protocol implementation.
Moreover, the performance of LLMs may be less effective for domain-specific tasks that require specialized knowledge, primarily due to their limited exposure to relevant data. To address this, we employ the background-augmented prompting technique [51], which leverages domain knowledge to construct task-specific prompts. Specifically, given the state description $s$ and the downstream task instruction $i$, e.g., ”summarize all state transitions”, we concatenate $s$ and $i$ to form a prompt that drives the augmented LLM to infer the state transitions.
@@ IV-B2 Message Types & States
The prompt templates employed to extract the set of message types and states are illustrated in Figure LABEL:fig:prompt_mes_sta. By incorporating previously extracted code paths, these template ensure that the LLM maintains contextual continuity. These prompts guide the augmented LLM in analyzing the previously obtained code files, thereby facilitating the extraction of message types and states defined within the protocol implementation. Fundamentally, our approach leverages the LLM’s code analysis capabilities to extract states and message types from structures or enumerated type variables. It is worth noting that our method can infer the FSM from the protocol implementation only when explicit definitions of states and message types are present. Nevertheless, our findings indicate that the majority of protocol implementations provide such explicit definitions, thereby making our approach viable.
Additionally, we specify the Desired Format in the prompt to ensure that the LLM outputs machine-readable results. The Augmented LLM will produce all message types and states in JSON format. These message types and states, encapsulated in the JSON data, will then be utilized to assist the augmented LLM in inferring state transitions.
@@ IV-B3 State Transitions
The prompt template employed by  ProtocolGPT to infer state transitions is depicted in Figure 6. We guide the augmented LLM to infer state transitions and the corresponding message types by utilizing code paths, predefined states, and message types. Inferring all state transitions within a protocol implementation is inherently a highly complex task. If the LLM were to attempt this task in a single step, it may result in suboptimal performance. Instead, we guide the LLM to systematically infer state transition relationships by iterating through all states.
Machine-readable FSM. We also provide Desired format in the prompt, which defines the pattern for state transitions. The LLM outputs state transitions in JSON format, following this specified pattern. Each JSON object includes a current_state key, whose value is an array of objects. Each object in the array contains two properties: receive_message and next_state. The current_state represents the peer’s current status, the next_state denotes the adjacent state to which the transition occurs, and the receive_message specifies the type of message triggering the state transition. The peers can directly transition from the current_state to the next_state without passing through any intermediate states. The example provided in Figure 1 illustrates the state transition. If the current_state is WAIT_CERT and a certificate message is received, the server transitions to the WAIT_CV state. Conversely, if an empty certificate message is received, the server transitions to the WAIT_FINISHED state. To capture all possible state transitions, ProtocolGPT generates multiple prompts based on the extracted states and the template shown in  Figure 6, systematically replacing the current_state field with previously identified states until all states are traversed.
Mitigation of the LLM Hallucination LLMs often generate plausible but non-factual predictions when applied to domain-specific tasks, referred as the hallucination of LLMs[22]. To mitigate the influence of the LLM’s stochastic behavior, we conduct 20 iterations of dialogue using the augmented model to extract states, message types, and state transitions. We then select the responses that appear with a probability greater than 80% across all iterations as the final result.
@@ IV-C Implementation
We implemented our proposed method using the LangChain framework[55] which is a widely adopted open-source framework for developing agents based on LLMs. The framework provides an extensive suite of interfaces, including Model I/O, Retrieval, and Vector Stores, allowing developers to implement customized functionalities.  ProtocolGPT implements both LLM augmentation and FSM inference functionalities within the LangChain framework. During the LLM augmentation phase, regular expressions are first used to match and filter out code files which are unrelated to the FSM. Documents are then segmented into manageable code chunks by our proposed syntax-aware splitting strategy which is implemented based on the text splitter provided by LangChain. The code chunks are then embedded by the model provided by OpenAI[53] and FAISS[54] and stored in a vector store. Ultimately  ProtocolGPT employs the chain-of-thought technique and background-augmented prompting technique to guide the augmented LLM in inferring the protocol state machine. We will release the source code of  ProtocolGPT and experimental data.

### v4

@@ IV System Design of ProtocolGPT
Motivated by the findings in Section III, we propose an approach (ProtocolGPT) for inferring state machines based on LLM. The core idea of our approach is to harness the code analysis capabilities of LLM to infer protocol state machines from protocol implementations. To overcome the challenges aforementioned, we use the augmented LLM that combines the strengths of both retrieval models and generative models. As shown in Figure 1, ProtocolGPT consists of two components: LLM augmentation and state machine inference.
@@
LLM Augmentation with Source Code. Initially, we isolate the FSM-related module from the complicated protocol implementations. To guarantee efficient statement inference, the source code is segmented into smaller chunks to bypass limitations of LLM’s context window. Moreover, ProtocolGPT maintains the integrity of the original code structure by syntax-aware segmenting for classes, functions, and other structural components. Our method preserves the semantic and contextual information, mitigating potential losses by traditional text segmentation technique. By embedding the LLM with these code chunks,  ProtocolGPT could more precisely retrieve the most relevant code during the state machine inference.
State Machine Inference. To enable the augmented LLM to infer FSM, we integrate the chain-of-thought [15] and background-augmented prompting techniques [16]. Leveraging domain-specific expertise, we formulate prompts tailored to the task at hand, which guide the augmented LLM in retrieving code segments that define both messages and states. The augmented LLM subsequently generates the message types and states based on the retrieved code. Finally, using the extracted message types and states, we traverse the entire state space to infer the corresponding transitions.
@@ IV-A LLM Augmentation with Source Code
While LLMs exhibit remarkable proficiency in code analysis, their efficacy diminishes significantly when tasked with analyzing large-scale programs. To address this limitation, ProtocolGPT leverages retrieval-augmented generation (RAG) techniques[32], which empower LLMs by integrating customizable external knowledge sources, such as code repositories. Although RAG models mitigate the limitation of relying on the training data, their performance deteriorates significantly if the external knowledge is not preprocessed. Specifically, protocol implementations often exceed the LLM’s context window or include irrelevant information, leading to side effects on retrieval. Consequently, preprocessing the code is essential for LLM augmentation.
@@ IV-A1 Code Filtering
To ensure scalability and maintainability, protocol implementations are typically structured into distinct modules, including state machines, configuration, testing, etc. Filtering out irrelevant cod not only reduces the cost of LLM prompts but also enhances the accuracy and efficiency of retrieval. Moreover, the tokens in protocol implementations significantly exceeds the context window of LLM. Therefore, extracting state machine module from the codebase forms the first step of the LLM augmentation.
State machine Module Identification. We observe that variables, constants, structures, and functions related to the state machine are often named according to terms defined in the RFCs. For instance, the CLIENT_HELLO struct in s2n-tls[28] corresponds to the client hello message in the RFC. Additionally, keywords such as state and message frequently appear in these context-specific names. Code filter identifies the directory containing the state machine module within the protocol implementation through regular expression matching. Based on these observation, we define a comprehensive set of keywords related to protocol state machines, derived from both RFCs and expert knowledge. The keyword set comprises the message types and states as defined in the RFC, supplemented by custom keywords based on expert knowledge. For instance, in the IKE protocol, Security Association (SA) appears frequently since the state of the SA essentially represents the state of the peer. These specially keywords facilitate the identification of the code related to state machine. Furthermore, our keyword set can be extended to accommodate additional protocols with negligible engineering effort, as these keywords can easily be found in the RFC section headings.
This pre-constructed keyword set is employed to match regular expressions in each document of the protocol implementation, and matching documents are marked as state machine-related code. The subdirectory where the documents have the highest match rate is chosen as the state machine module, providing domain knowledge for LLM augmentation.
@@ IV-A2 Syntax-aware Code Segmentation
For the RAG model, the retrieved data is directly provided as context to the prompt, enabling the LLM to generate a response. To retrieve data from the vector database, the documents are transformed into chunks and embeded by the embedding model. However, traditional character-based splitting strategies risk disrupting semantically meaningful elements and may result in the loss of contextual information. Consequently, two critical factors must be taken into account during code segmentation: the chunk size and the syntactic structure of the code.
@@
Appropriate Chunk Size Adjustment. The input provided to the LLM includes not only the retrieved chunks but also instructions. Appropriately sized chunks contain more precise contextual information, resulting in improved matching accuracy. However, if the length of the retrieved chunks approaches or exceeds the context window of LLM, the performance of the model is adversely affected.
@@
Syntactic Program Structure Construction. Conventional text segmentation overlooks the structural composition of the document. When applied to source code, which is characterized by inherent syntax, these strategies can break apart key components like the classes and functions. Moreover, the rich contextual information embedded within the code may be degraded during the segmentation process.
Figure 2: Workflow of Syntax-aware Code Segmentation.
@@
Building upon the preceding observations, we propose a syntax-aware code splitting strategy. This method breaks down code into appropriately sized chunks, ensuring that the syntactic integrity of the code remains intact while also preserving a sufficient level of contextual information. Figure 2 illustrates the workflow of syntax-aware code segmentation. (1) Initially, we define a maximum chunk size, referred to as MaxChunkSize. We then recursively decompose the documents using a series of language-specific separators until the size of each chunk is smaller than MaxChunkSize. These separators consist of keywords that define critical elements such as structures, classes, functions, and other constructs. For instance, in the C programming language, separators include terms such as struct, class, enum, etc. To facilitate the application of our method across multiple programming languages, we compile a comprehensive list of separators for C, C++, Python, and others. (2) To prevent the creation of overly small chunks that lack sufficient context, we establish a predefined minimum threshold for chunk size, named MinChunkSize. If the length of the chunk falls below MinChunkSize, it is merged with an adjacent chunk to ensure coherence. (3) Finally, to preserve contextual continuity between chunks, we regenerate new chunks based on previously segmented ones, ensuring a specified degree of overlap is maintained between them.
@@ IV-A3 Vector Store
The augmented model employs semantic search to retrieve relevant code from a vector store. These code snippets are then appended to the original prompt, which is subsequently input into the generative model. Retrieval engine is facilitated by converting code into embeddings (high-dimensional vectors). The vector store is specifically optimized for storing embeddings of code, enabling the use of advanced search algorithm to identify matches between similar vectors.
When ProtocolGPT inferring state machine , it initially embeds the prompt to create an embedded version. It then retrieves the vectors from the vector store that most similar to the embedded prompt. Finally, the augmented LLM generate a more contextually relevant response based on the original prompt and the code corresponding to these retrieved vectors.
For the storage and retrieval of unstructured code, the optimal approach is to embed it and store the resulting vectors. Our approach leverages an embedding model provided by OpenAI[33] to transform code into vector data. The vector data is subsequently stored and retrieved using a retrieval engine facilitated by FAISS[34]. Each vector corresponds to a specific piece of code chunk. The retrieval engine employs the approximate nearest neighbor algorithm to retrieve vectors relevant to the prompt.
@@ IV-B FSM Inference
Before inferring FSM, it is essential to define the FSM that the model generates. Our primary objective is to examine the transition relationships between different states within the implementation of the protocol. Consequently, we anticipate the FSM generated by the LLM to serve as a high-level model that encompasses both the states and the transitions between these states. Specifically, the FSM details the potential states the protocol server can attain, the pairs of source and destination states, and the conditions necessary for state transitions. To facilitate support for downstream applications, FSM discussed in this paper is assumed to be a non-deterministic FSM that can be transformed into a deterministic FSM. Notably, an non-deterministic FSM may possess multiple initial states, and a single state can transition to several successor states in response to diverse inputs. Inspired by RFCNLP[6] and StateLifter[7], we define an non-deterministic FSM in Definition 1 as a formal model.
@@
The FSM inference phase leverages prompt engineering to guide the augmented LLM in deriving FSM from the implementations. However, this process is not instantaneous. As task complexity increases with more steps, the LLM’s probability of errors also rises. In contrast, the model demonstrates exceptional accuracy when executing single instructions. To mitigate this limitation, we employ the COT  [15] to guide the augmented LLM. Specifically, the FSM inference is structured into three stages: (1) Retrieving the code snippets that define states, messages, and state transitions. (2) Extracting all the messages and states defined in the protocol implementation. (3) Identifying all the state transitions and the corresponding message types that trigger these transitions. By following these steps,  ProtocolGPT gradually extracts the information which forms the FSM of the protocol implementation.
Moreover, the performance of LLMs may be less effective for domain-specific tasks, primarily due to their limited exposure to relevant data. To address this, we employ the BAP [16], which leverages domain knowledge to construct task-specific prompts. Specifically, given the state description $s$ and the downstream task instruction $i$, e.g., ”summarize all state transitions”, we concatenate $s$ and $i$ to form a prompt that drives the augmented LLM to infer the state transitions.
@@ IV-B2 Message Types & States
The prompt templates employed to extract the set of message types and states are illustrated in Figure LABEL:fig:prompt_mes_sta. By incorporating previously extracted code paths, these template ensure that the LLM maintains contextual continuity. These prompts guide the augmented LLM in analyzing the previously obtained code files, thereby facilitating the extraction of message types and states defined within the protocol implementation. Fundamentally, our approach leverages the LLM’s code analysis capabilities to extract states and message types from structures or enumerated type variables. It should be noted that our method can infer the FSM from the protocol implementation only when explicit definitions of states and message types are present. Nevertheless, our findings indicate that the Top-50 most widely used open-source protocol implementations provide such explicit definitions, making our approach viable.
@@ IV-B3 State Transitions
The prompt template employed by  ProtocolGPT to infer state transitions is depicted in Figure 5. We guide the augmented LLM to infer state transitions and the corresponding message types by utilizing code paths, predefined states, and message types. Inferring all state transitions within a protocol implementation is inherently a highly complex task. If the LLM were to attempt this task in a single step, it may result in suboptimal performance. Instead, we guide the LLM to systematically infer state transition relationships by iterating through all states. Specifically, ProtocolGPT generates multiple prompts using the extracted states. Then it iteratively replaces the current_state with all previously identified states.
Machine-readable FSM. A significant challenge in utilizing LLMs for downstream tasks is ensuring that their outputs are machine-readable. Unlike static and dynamic analysis methods, the outputs of generative model are formulated in natural language. Therefore, without imposing constraints on the output of model, the resulting FSM would be unusable. This challenge can be addressed through fine-tuning prompts, which guide the model to generate outputs according to the given patterns. Specifically, we provide desired format in the prompts, which defines the pattern for states, message types, and state transitions. The augmented LLM generates FSM in JSON format, following this specified pattern. For example, each transition includes a current_state key, whose value is an array in  Figure 5. Each object in the array contains two properties: receive_message and next_state. The current_state represents the peer’s current state, the next_state denotes the destination state, and the receive_message specifies the type of message triggering the transition.
Mitigation of the LLM Hallucination. LLMs often generate plausible but non-factual predictions when applied to domain-specific tasks, referred as the hallucination of LLMs[14]. To mitigate stochastic behavior of the LLM, we conduct 20 iterations of dialogue using the augmented model to extract states, message types, and state transitions. We then select the responses that appear with a probability greater than 80% across all iterations as the final result.
@@ IV-C Implementation
We implement our method with LangChain framework which is a widely adopted open source framework to develop agents based on LLM. The framework provides an extensive suite of interfaces, including Model I/O, Retrieval, and Vector Stores, allowing developers to implement customized functionalities.  ProtocolGPT implements both LLM augmentation and FSM inference functionalities within the LangChain. During the LLM augmentation, regular expressions are first used to extract state machine module. Documents are then segmented into manageable code chunks by our proposed syntax-aware code segmentation strategy which is implemented based on the text splitter provided by LangChain. The code chunks are then embedded by the model provided by OpenAI and FAISS and stored in a vector store. Ultimately  ProtocolGPT employs the COT and BAP to guide the augmented LLM to infer FSM. The source code and the experimental data will be released immediately after the work is accepted.